OpenStudy (anonymous):

can someone tell me if this is correct write the slope intercept equation for the line with a slope 1/3 and y-intercept (0,-6) y=1/3(0)-6 x=0,y=-6 And, is this correct? write the slope intercept equation for the line that passes through (-7,-1) and has a slope -6/7 y=mx+b y=-6/7/(-7)/1)-1 y=-42/7-1 y=-6+-1 x=-6,y=-1 IS THIS ? CORRECT

OpenStudy (anonymous):

y = 1/3x -6

OpenStudy (anonymous):

o k

OpenStudy (lalaly):

\[y--6 = 1/3 (x-0)\]\[y=1/3 x -6\]

OpenStudy (lalaly):

\[y+1=-6/7 x -7\]\[y--1 = -6/7 (x--7)\]

OpenStudy (lalaly):

\[y=-6/7 x-5\]

OpenStudy (anonymous):

You are given a point and a slope, so you need to start with the point-slope formula, then work it into a slope intercept form.

OpenStudy (lalaly):

\[y-y1 = m(x-x1)\]

OpenStudy (anonymous):

and this is the answer?